OpenStudy (anonymous):

aaron kicked a soccer ball with an initial velocity of 39 feet per second at an angle of 44 degrees with the horizontal after 0.9 second, how far has the ball traveled horozontally?

OpenStudy (anonymous):

26. 25.245?

OpenStudy (anonymous):

27. 6.081

OpenStudy (perl):



OpenStudy (perl):

you can make parametric equations for the flight $$x(t) = 39 \cos(44 \deg ) t\\ y(t) = 39 \sin(44 \deg) t - \frac{1}{2} g~t^2$$

OpenStudy (perl):

$$ \Large x(t) = 39 \cos(44 ^\circ ) t\\ \Large y(t) = 39 \sin(44^\circ ) t - \frac{1}{2} g~t^2 $$

OpenStudy (perl):

so you can answer the first question, find x (0.9)

OpenStudy (anonymous):

25.25?

OpenStudy (perl):

After 0.9 second, how far has the ball traveled horizontally? $$ \Large x(t) = 39 \cos(44 ^\circ ) t\\ \Large x(0.9) = 39 \cos(44 ^\circ )~\cdot (0.9)\\ $$

OpenStudy (perl):

yes thats correct